OpenStudy (anonymous):

Use the graph of f(x) = log x to obtain the graph of g(x) = log x + 5. I will post the graphs

OpenStudy (anonymous):

@rafabc02

OpenStudy (mathteacher1729):

This problem is PERFECT for Geogebra. http://www.geogebra.org/cms/installers It's free, works on Windows & Mac. :)

OpenStudy (anonymous):

You can evaluate at x=0 and see what you get. Then you shoud remmeber the graph of log (x).

OpenStudy (anonymous):

Sorry, I mean evaluate at x=1

OpenStudy (anonymous):

i dont understand :/

OpenStudy (anonymous):

thnx for the tip @mathteacher1729

OpenStudy (anonymous):

have u studied graph transformations?

OpenStudy (anonymous):

i did for quadratic functions

OpenStudy (anonymous):

then do u know what happens in this case: say, f(x) = x^2 and g(x) = x^2 + 1 then how will u transform f(x) to get g(x) ?

OpenStudy (anonymous):

g(x) is transformed up 1? or was it left 1?

OpenStudy (anonymous):

its up 1. because the transformation is happening to the y-variable. so the graph will move along y-axis. for a better explanation, view this page: http://www.purplemath.com/modules/fcntrans.htm so, now can u relate this with ur question above?

OpenStudy (anonymous):

so its B? but when i plug it into my calculator i get D..

OpenStudy (anonymous):

yes, it should be B. i dont know why u're getting D.

OpenStudy (anonymous):

ok thnx!

OpenStudy (anonymous):

hang on!

OpenStudy (anonymous):

yes, its B for sure!

OpenStudy (anonymous):

thank u soo much!

OpenStudy (anonymous):

sure, no problem! :)

OpenStudy (anonymous):

go through the link i gave before (about the graph transformation) and u'll be able to solve such problems without much difficulty!